# FALCON: Feedback-driven Adaptive Long/short-term memory reinforced Coding OptimizatioN

### v3

@@ Abstract
Recently, large language models (LLMs) have achieved significant progress in automated code generation. Despite their strong instruction-following capabilities, these models frequently struggled to align with user intent in the coding scenario. In particular, they were hampered by datasets that lacked diversity and failed to address specialized tasks or edge cases. Furthermore, challenges in supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) led to failures in generating precise, human-intent-aligned code. To tackle these challenges and improve the code generation performance for automated programming systems, we propose Feedback-driven Adaptive Long/short-term memory reinforced Coding OptimizatioN (i.e., FALCON). FALCON is structured into two hierarchical levels, from the global level, long-term memory improves code quality by retaining and applying learned knowledge, while from the local level, short-term memory allows for the incorporation of immediate feedback from compilers and AI systems. Additionally, we introduce meta-reinforcement learning with feedback rewards to solve the global-local bi-level optimization problem and enhance the model’s adaptability across diverse code generation tasks. Extensive experiments are conducted and it is found that our technique achieves state-of-the-art performance, leading other reinforcement learning methods by more than 4.5 percentage points on the MBPP benchmark and 6.1 percentage points on the Humaneval benchmark. The open-sourced code is publicly available at https://anonymous.4open.science/r/FALCON-3B64/README.md.
@@ I Introduction
The development of Large Language Models (LLMs) has significantly advanced automated code generation [1]. Models like CodeLLaMA [2] and DeepSeek-Coder [3], tailored for code-centric tasks, have demonstrated outstanding performance across programming challenges. While LLMs excel in instruction-following through tuning [4], they often misalign with user intent, making feedback-based adjustments critical. For example, InstructGPT [5] leverages reinforcement learning with human feedback (RLHF), and CodeRL [6] uses compilation feedback to refine model performance. Similarly, CompCoder [7] enhances code compilability with compiler feedback, and RLTF [8] offers fine-grained feedback on compiler errors. However, current RL frameworks generate compilation errors and overlook non-differentiable features (e.g. coding style) that affect the performance significantly [4]. To address these challenges, we propose a reinforcement learning system combining long-term and short-term memory feedback. From the global level, long-term memory tracks trends over time for higher-quality code retrieval, while from the local level, short-term memory captures recent errors and immediate feedback. The main contributions of this paper are as follows:
@@
  • Meta-Reinforcement Learning for Generalization Across Tasks: We enhance the model’s versatility by incorporating meta-reinforcement learning, allowing it to generalize across diverse programming tasks efficiently, adapt to new environments quickly, and handle a wide range of coding challenges with fewer training iterations.
@@ II-A Pre-trained Models for Code Generation
In recent years, pre-trained language models have made significant progress in the field of code generation. Trained on large-scale code corpora, these models have demonstrated powerful code generation capabilities. For example, CodeBERT [9], a model based on an encoder-only architecture, has shown impressive results. With the advent of in-context learning methods, decoder-only Transformer models have become the dominant technology for language modeling [10]. Several models, such as CodeGPT [11], CodeGeeX [12], and DeepSeek-Coder [3], use Causal Language Modelling (CLM) pretraining, while others like CodeT5 [13] and AlphaCode [14] utilize an encoder-decoder architecture. Additionally, models like CodeFusion [15] leverage diffusion-based techniques for code generation. These pre-trained models exhibit great potential in code generation tasks, achieving notable improvements in accuracy through various architectures and training strategies. However, they still face challenges in ensuring the syntactical and functional correctness of the generated code.
@@ II-B Reinforcement Learning on code
Reinforcement learning (RL) is a method that learns optimal strategies through reward signals from interacting with the environment [16]. It excels in sequence generation tasks, such as enhancing performance in translation and summarization models by improving BLEU and ROUGE scores [17]. Unlike traditional natural language processing tasks, code generation is more complex. Beyond syntactical correctness, the generated code must be functionally accurate, meaning it should compile and perform as expected in various scenarios. Passing unit tests is important for verifying correctness, safety, and precision, but not sufficient unless the tests are comprehensive and well-designed. Unit tests must cover diverse use cases and edge cases to ensure the code adheres to standards and meets requirements. For instance, the RL-based code generation fine-tuning framework, CodeRL [6], guides fine-tuning by integrating unit test signals with reinforcement learning. PPOCoder [18] improves on CodeRL by employing Proximal Policy Optimization to refine the approach, while RLTF [8] incorporates specific error information from the code and multi-granularity feedback alongside an online framework for model training. StepCoder [19] enhances the code generation process through compiler feedback and segmental optimization. However, current RL-based methods still face some limitations. Primarily, these methods lack detailed feedback on specific types and distributions of programming errors, which is crucial for identifying patterns, understanding root causes, and creating targeted interventions. This makes addressing recurring issues difficult. Additionally, the corrective mechanisms are not robust enough, often failing to provide timely, specific guidance, hindering effective learning. Additionally, the diversity of input-output examples in existing benchmarks is limited, restricting the model’s ability to adapt and generalize to different or unseen problems [8].
@@ III Problem Setting
We aim to enhance the automated code generation capabilities of Large Language Models (LLMs) by addressing key challenges in accuracy, diversity, error correction, and code quality. Formally, given a high-level specification $D$, the task is to generate a sequence of code tokens $W=\{w_{1},w_{2},\ldots,w_{T}\}$ that maximizes the conditional probability $P(W|D,\theta)$, where $\theta$ represents the model parameters. The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}\left[\log P(W|D,\theta)\right]$.
Challenges. Effective code generation is impeded by several factors. Accuracy requires minimizing syntactical or logical errors in $W$ to ensure correct functionality. Diversity of input-output examples $\mathcal{X}$ is often limited, restricting the model’s ability to generalize across varied programming tasks. Efficient error correction mechanisms are necessary to identify and rectify errors in $W$, ensuring robust performance. Maintaining high code quality, which encompasses adherence to coding standards, style guidelines, and managing code complexity $C(W)$, remains a persistent challenge. Additionally, adaptability refers to the model’s capacity to adapt to new tasks and incorporate feedback for continuous improvement, which is constrained without robust memory mechanisms.
FALCON Framework. To address these challenges, we propose the FALCON framework. The framework leverages both a long-term memory buffer $\mathcal{M}_{\text{long}}=\{(D_{i},W_{i},T_{i},F_{i})\}_{i=1}^{N}$ and a short-term memory buffer $\mathcal{M}_{\text{short}}=\{(D_{j},W_{j},T_{j},F_{j})\}_{j=1}^{M}$ to utilize diverse feedback, enabling fine-tuning of $\theta$. Formally, FALCON seeks to optimize $\theta$ by maximizing a composite reward function $R(W,F)=\alpha T(W)+\beta S(W)+\gamma C(W)+\delta E(W)$, where $T(W)$, $S(W)$, $C(W)$, and $E(W)$ represent unit test, code style, complexity, and error feedback, respectively. The optimization objective is defined as $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}\left[\mathbb{E}_{W\sim P(W|D,\theta)}\left[R(W,F)\right]\right]$.
Assumptions. The effectiveness of FALCON is based on the following assumptions: 1. Exchangeability of the dataset $\mathcal{D}=\{D_{i}\}_{i=1}^{N}$, implying that the order of tasks does not influence model performance; 2. Independence of feedback signals $F$ given the generated code $W$; 3. Sufficient memory capacity in both $\mathcal{M}_{\text{long}}$ and $\mathcal{M}_{\text{short}}$ to store relevant interactions without significant data loss; and 4. Feedback efficacy, ensuring that integrated feedback mechanisms provide meaningful and actionable information to guide the optimization of $\theta$.
Optimization Objective. Given a set of tasks $\mathcal{T}=\{D_{i}\}_{i=1}^{N}$, the goal is to learn optimal parameters $\theta^{*}$ that maximize the expected composite reward across all tasks, formulated as $\theta^{*}=\arg\max_{\theta}\frac{1}{N}\sum_{i=1}^{N}R(W_{i},F_{i})$, where $W_{i}\sim P(W|D_{i},\theta)$ is the generated code for task $D_{i}$, and $F_{i}$ represents the aggregated feedback. The formulation ensures continuous improvement by leveraging both historical and recent feedback to enhance code generation quality.
@@ IV Methodology
In this section, we explore the FALCON framework, which integrates comprehensive unit testing with reinforcement learning, supported by both long-term and short-term memory buffers. During the code generation process, the system stores task descriptions, generated code, and various feedback (e.g., compilation results, code style, and complexity) in the long-term memory buffer. By retrieving this information, the model references high-quality code, avoids past mistakes, and ensures adherence to required standards. After generating the code, a judge model evaluates it and calculates rewards based on the feedback, which are then used to update the model’s parameters through reinforcement learning. All generated code and feedback are stored for future reference and optimization. The combination of long-term and short-term memory feedback in the FALCON framework allows the model to not only learn from a wide range of historical data but also adapt quickly to new tasks based on recent performance. The overall framework is illustrated in Figure 1.
Figure 1: The overview of the FALCON framework. Detailed feedback process (see Appendix Figures 5 and 6 for feedback cases). Reinforcement learning updates using feedback to compute RL loss and optimize the model (see Figure 2 for the detailed meta-reinforcement learning framework).
IV-A Task Definition: Code Generation by LLMs
The code generation task involves the automated creation of computer code $W$ from a high-level specification of the desired behavior $D$. The goal of this process is to enhance the likelihood of generating code that effectively solves a given problem, which is represented as a conditional probability $P(W|D)$. This task can be modeled as a sequence-to-sequence problem, aiming to maximize the conditional probability of generating the correct output given the input and the parameters of the LLM model, mathematically represented as:
$\max_{\theta}P(W|D,\theta)=\max_{\theta}\prod_{t=1}^{T}p(w_{t}|D,\theta,w_{1:t-1})$ (1)
where $\theta$ denotes the parameters of the LLM model. The LLM model is employed to learn the conditional distribution and is trained on a set of input-output pairs. During the training phase, the parameters $\theta$ are optimized to increase the likelihood of generating accurate outputs for a given input.
IV-B MRLF: Meta-Reinforcement Learning with Dual Memory Buffers
@@
IV-C Long-Term Memory Feedback
Given that retrieving information from long-term memory helps improve code quality, we use the FAISS framework [22] to retrieve relevant historical code, feedback, and evaluation scores. Task descriptions and feedback are transformed into embedding vectors and then indexed. During code generation, a query vector from the current task retrieves the top-k most similar historical data to guide the process and avoid past errors. The prompt template is provided in the appendix. Consider a set of historical data $\mathcal{D}={(t_{i},f_{i},e_{i})}_{i=1}^{n}$, where $t_{i}$ represents the task description, $f_{i}$ is the corresponding feedback, and $e_{i}$ is the evaluation score. We use an embedding function $\phi(\cdot)$ to transform these tasks and feedback into embedding vectors $\boldsymbol{v}_{i}=\phi(t_{i},f_{i})$ and index them with FAISS.
During the code generation phase, the current task description $\boldsymbol{t}_{\text{current}}$ and feedback $\boldsymbol{f}$ are transformed into a query vector $\boldsymbol{q}=\phi(\boldsymbol{t}_{\text{current}})$. We compute the similarity between the query vector $\boldsymbol{q}$ and the historical vectors $\boldsymbol{v}_{i}$ using cosine similarity $\text{cos}(\boldsymbol{q},\boldsymbol{v}_{i})$, and retrieve the top-$k$ most similar historical tasks. The retrieval process can be represented as:
IV-D Short-Term Memory Feedback
During the reinforcement learning phase, we utilize the generated code $\hat{w}$ to construct the reinforcement learning loss function as follows:
$L_{r1}=-\sum_{t=S_{\text{fine}}}^{E_{\text{fine}}}R_{\text{fine}}(\hat{w}_{t})\log p(\hat{w}_{t}|D,\theta,\hat{w}_{1:t-1})$ (8)
where $R_{\text{fine}}(\ast)$ represents the reward coefficient, and $S_{\text{fine}}$ and $E_{\text{fine}}$ denote the start and end positions of the code snippet, respectively. These values are determined based on different types of feedback.
where $R_{\text{coarse}}$ is based on compiler feedback with the start position $S_{\text{coarse}}=0$ and end position $E_{\text{coarse}}=T$.
Coding Style Feedback. To further enhance the quality of the generated code, we employ AI Feedback to optimize coding style. An evaluation model scores the generated code based on adherence to the expected coding style standards. The scoring system ranges from -1 to 2, and these evaluation scores are directly used as reward signals in the reinforcement learning process to guide the model toward producing higher-quality code. The coding style assessment template is provided in Table X.
Complexity Feedback. Similar to the coding style, we use AI Feedback to evaluate complexity and calculate rewards based on the scores. The complexity assessment template is provided in Table XI.
Long-term Error Type Feedback. Introducing a reward mechanism that combines short-term memory recall error rate, current test error, and long-term memory feedback of past task performance enables the model to dynamically adjust its rewards based on past error patterns, adapt to various error types and feedback, and ultimately enhance its generalization ability:
where $N_{\text{error}}$ represents the frequency of each error type in the generated code, which is the short-term memory feedback and the immediate feedback for the current task. $P_{\text{error}}$ represents the proportion of each error type, which is the long-term memory feedback. This is the model’s performance statistics on long-term tasks, providing a cumulative history of various error types. By correct rewarding, the model can reduce the occurrence of these errors and thereby enhance the accuracy and quality of the generated code.
@@ V-B Quantitative Evaluation on HumanEval and MBPP
To further validate the effectiveness of our method, we evaluate the zero-shot performance of the DeepSeek-Coder-Instruct model trained with our method on APPS, using the well-established MBPP and HumanEval benchmarks. We also compare these results against other reinforcement learning methods, such as PPOCoder and RLTF. The experimental results are illustrated in Table II.
@@
TABLE III: Test set results of AGENTBENCH.
Model Size VER OS DB KG
GPT-3.5-turbo - 0613 31.6 15.7 25.9
GPT-4 - 0613 42.4 32.0 58.8
Tinyllama 1.1B - 2.8 0 0
Codellama 7B instruct 9.7 2.7 0
Qwen 7B chat 12.5 13.0 7.0
Agentlm 7B chat 14.6 33 9.0
Deepseek-Coder 6.7B instruct 17.4 23.3 6.8
DeepSeek-Coder (FALCON) 6.7B instruct 22.2 26.7 9.0

### v5

@@ Abstract
Recently, large language models (LLMs) have achieved significant progress in automated code generation. Despite their strong instruction-following capabilities, these models frequently struggled to align with user intent in the coding scenario. In particular, they were hampered by datasets that lacked diversity and failed to address specialized tasks or edge cases. Furthermore, challenges in supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) led to failures in generating precise, human-intent-aligned code. To tackle these challenges and improve the code generation performance for automated programming systems, we propose Feedback-driven Adaptive Long/short-term memory reinforced Coding OptimizatioN (i.e., FALCON). FALCON leverages long-term memory to retain and apply learned knowledge, short-term memory to incorporate immediate feedback, and meta-reinforcement learning with feedback rewards to address global-local bi-level optimization and enhance adaptability across diverse code generation tasks.Extensive experiments show that FALCON achieves state-of-the-art performance, outperforming other reinforcement learning methods by over 4.5% on MBPP and 6.1% on Humaneval, with the code publicly available.https://anonymous.4open.science/r/FALCON-3B64/README.md.
@@
Figure 1: The overview of the FALCON framework. Detailed feedback process. Reinforcement learning updates using feedback to compute RL loss and optimize the model (see Figure 2 for the detailed meta-reinforcement learning framework).
@@ I Introduction
The development of Large Language Models (LLMs) has significantly advanced automated code generation [1]. Models like CodeLLaMA [2] and DeepSeek-Coder [3], tailored for code-centric tasks, have demonstrated outstanding performance across programming challenges. While LLMs excel in instruction-following through tuning [4], they often misalign with user intent, making feedback-based adjustments critical. For example, InstructGPT [5] leverages reinforcement learning with human feedback (RLHF), and CodeRL [6] uses compilation feedback to refine model performance. Similarly, CompCoder [7] enhances code compilability with compiler feedback, and RLTF [8] offers fine-grained feedback on compiler errors. However, current RL frameworks generate compilation errors and overlook non-differentiable features (e.g. coding style) that affect the performance significantly [4]. To address these challenges, we propose a reinforcement learning system combining long-term and short-term memory feedback. From the global level, long-term memory tracks trends over time for higher-quality code retrieval, while from the local level, short-term memory captures recent errors and immediate feedback. In summary, the main contributions of this paper are as follows:
@@
  • Meta-Reinforcement Learning for Generalization Across Tasks: We enhance the model’s versatility by incorporating meta-reinforcement learning, allowing it to efficiently generalize across diverse programming tasks, adapt quickly to new environments, and handle a wide range of coding challenges with fewer training iterations.
@@ II-A Pre-trained Models for Code Generation
In recent years, pre-trained language models have made significant progress in the field of code generation. Trained on large-scale code corpora, these models have demonstrated powerful code generation capabilities. For example, CodeBERT [9], a model based on an encoder-only architecture, has shown impressive results. With the advent of in-context learning methods, decoder-only Transformer models have become the dominant technology for language modeling [10]. Several models, such as CodeGeeX [11] and DeepSeek-Coder [3], use Causal Language Modelling (CLM) pretraining, while others like CodeT5 [12] utilize an encoder-decoder architecture. Additionally, models like CodeFusion [13] leverage diffusion-based techniques for code generation. These pre-trained models exhibit great potential in code generation tasks, achieving notable improvements in accuracy through various architectures and training strategies. However, they still face challenges in ensuring the syntactical and functional correctness of the generated code.
@@ II-B Reinforcement Learning on code
Reinforcement learning (RL) learns optimal strategies through interaction with the environment and reward signals [14], showing strong performance in sequence generation tasks. Code generation, however, is more complex than typical NLP tasks, as it requires not only syntactic correctness but also functional accuracy—ensuring the code compiles and behaves as expected. While unit tests are helpful for verifying correctness, they are insufficient unless well-designed to cover diverse and edge cases. To address this, methods like CodeRL [6] incorporate unit test feedback into fine-tuning via reinforcement learning. PPOCoder [15] improves this with Proximal Policy Optimization, and RLTF [8] further integrates error information and multi-granularity feedback. StepCoder [16] enhances the process using compiler feedback and segmental optimization. Despite these advances, existing RL-based approaches often lack fine-grained feedback on error types and distributions, limiting their ability to address recurring issues. Moreover, their corrective mechanisms are typically weak, offering limited and untimely guidance. Finally, the limited diversity in benchmark datasets restricts model generalization to unseen scenarios [8].
@@ III Problem Setting
We aim to enhance automated code generation capabilities of Large Language Models (LLMs) by addressing accuracy, diversity, error correction, code quality, and adaptability. Formally, given a high-level specification $D$, the task is to generate a sequence of code tokens $W=\{w_{1},w_{2},\ldots,w_{T}\}$ that maximizes $P(W|D,\theta)$, with the objective $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}[\log P(W|D,\theta)]$. Accuracy requires minimizing syntactical or logical errors, while limited diversity in input-output pairs $\mathcal{X}$ restricts generalization. Efficient error correction is crucial for identifying and fixing errors in $W$. Code quality involves adhering to standards, style guidelines, and managing complexity $C(W)$. Adaptability depends on the model’s capacity to incorporate feedback for continuous improvement, which is constrained without robust memory. The FALCON framework leverages a long-term memory buffer $\mathcal{M}_{\text{long}}=\{(D_{i},W_{i},T_{i},F_{i})\}_{i=1}^{N}$ and a short-term memory buffer $\mathcal{M}_{\text{short}}=\{(D_{j},W_{j},T_{j},F_{j})\}_{j=1}^{M}$ to utilize diverse feedback for fine-tuning $\theta$. FALCON optimizes $\theta$ by maximizing a composite reward $R(W,F)=\alpha T(W)+\beta S(W)+\gamma C(W)+\delta E(W)$, where $T(W)$, $S(W)$, $C(W)$, and $E(W)$ denote unit test, style, complexity, and error feedback, respectively, yielding $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{D\sim\mathcal{D}}[\mathbb{E}_{W\sim P(W|D,\theta)}[R(W,F)]]$. We assume exchangeability of $\mathcal{D}$, independence of feedback $F$ given $W$, sufficient memory capacity in $\mathcal{M}_{\text{long}}$ and $\mathcal{M}_{\text{short}}$, and effective feedback mechanisms. Given tasks $\mathcal{T}=\{D_{i}\}_{i=1}^{N}$, the ultimate goal is $\theta^{*}=\arg\max_{\theta}\frac{1}{N}\sum_{i=1}^{N}R(W_{i},F_{i})$ with $W_{i}\sim P(W|D_{i},\theta)$, ensuring continuous improvement by leveraging both historical and recent feedback to enhance code generation quality.
@@ IV Methodology
The FALCON framework enhances code generation by LLMs through unit testing and reinforcement learning, leveraging long-term and short-term memory buffers. Task descriptions, generated code, and feedback (e.g., compilation results, code style, complexity) are stored in the long-term memory buffer, allowing the model to reference high-quality code, avoid errors, and adhere to standards. A judge model evaluates the code, calculates feedback-based rewards, and updates model parameters via reinforcement learning. This dual memory mechanism enables the model to improve code quality using historical data while adapting to new tasks. The method section includes three components: Meta-Reinforcement Learning with Dual Memory Buffers IV-A, introducing the dual memory structure; and Long-Term Memory Feedback IV-B, optimizing with historical data. See Figure 1.
IV-A MRLF: Meta-Reinforcement Learning with Dual Memory Buffers
@@
IV-B Long-Term Memory Feedback
Retrieving information from long-term memory helps improve code quality. We use the FAISS framework [19] to retrieve relevant historical code, feedback, and evaluation scores. Task descriptions and feedback are transformed into embedding vectors and then indexed. During code generation, a query vector from the current task retrieves the top-k most similar historical data to guide the process and avoid past errors. The prompt template is provided in the appendix. Consider a set of historical data $\mathcal{D}={(t_{i},f_{i},e_{i})}_{i=1}^{n}$, where $t_{i}$ represents the task description, $f_{i}$ is the corresponding feedback, and $e_{i}$ is the evaluation score. We use an embedding function $\phi(\cdot)$ to transform these tasks and feedback into embedding vectors $\boldsymbol{v}_{i}=\phi(t_{i},f_{i})$ and index them with FAISS. During the code generation phase, the current task description $\boldsymbol{t}_{\text{current}}$ and feedback $\boldsymbol{f}$ are transformed into a query vector $\boldsymbol{q}=\phi(\boldsymbol{t}_{\text{current}})$. We compute the similarity between the query vector $\boldsymbol{q}$ and the historical vectors $\boldsymbol{v}_{i}$ using cosine similarity $\text{cos}(\boldsymbol{q},\boldsymbol{v}_{i})$, and retrieve the top-$k$ most similar historical tasks. The retrieval process can be represented as:
IV-C Short-Term Memory Feedback
During the reinforcement learning phase, we utilize the generated code $\hat{w}$ to optimize the policy based on the reinforcement learning loss defined in Equation (1). As previously described, this loss incorporates reward signals such as unit test feedback, code complexity, and style norms, and is computed over the token span from $S_{\text{fine}}$ to $E_{\text{fine}}$.
$S_{\text{coarse}}=0,\quad E_{\text{coarse}}=T$
where $R_{\text{coarse}}$ is based on compiler feedback with the start and end positions set to 0 and $T$.
Coding Style Feedback. To further enhance the quality of the generated code, we employ AI Feedback to optimize coding style. An evaluation model scores the generated code based on adherence to the expected coding style standards. The scoring system ranges from -1 to 2, and these evaluation scores are directly used as reward signals in the reinforcement learning process to guide the model toward producing higher-quality code.
Complexity Feedback. Just like with coding style, we use AI Feedback to evaluate complexity and calculate rewards based on the scores.
Long-term Error Type Feedback. A reward mechanism combining short-term error recall, current test errors, and long-term memory of past performance dynamically adjusts rewards. The formula is:
where $N_{\text{error}}$ represents short-term feedback on current task errors, and $P_{\text{error}}$ represents long-term error proportions. This mechanism reduces errors and enhances the accuracy and quality of code generation.
@@ V-B Quantitative Evaluation on HumanEval and MBPP
To further validate the effectiveness of our method, we evaluate the zero-shot performance of the DeepSeek-Coder-Instruct model, trained with our method on our custom dataset, using the well-established MBPP and HumanEval benchmarks. We also compare these results against other reinforcement learning methods, such as PPOCoder and RLTF. The experimental results are illustrated in Table II.